The length of a rectangle is half the width. The area is 25 square meters. So whats the perimeter?
Answer by lwsshak3(11628)

The length of a rectangle is half the width. The area is 25 square meters. So whats the perimeter?
***
let x=length
2x=width
length*width=area
x*2x=25
2x^2=25
x^2=25/2
x=5/√2=5√2/2(length)
2x=5√2(width)
2(width+length)=perimeter
2(5√2+5√2/2)=10√2+5√2=15√2
perimeter=15√2 meters